# Advancing Model-Assisted Design in Phase I Trials through Bayesian Dose-Response Model

### v1

@@ 1 Introduction
Given these limitations, model-assisted designs (Zhou et al.,, 2018; Yuan et al.,, 2019), which merge the operational simplicity of algorithm-based designs with the good performance of model-based approaches, have attracted growing interest in recent years. In interval designs, dose escalation and de-escalation are determined based on the observed DLT rate at the current dose, and compared to a predefined toxicity tolerance interval to decide the appropriate dose for the next patient cohort. Oron et al., (2011) demonstrate that interval designs offer convergence properties similar to those of the CRM. Interval designs employ a model (e.g., the binomial model) to determine dose escalation and de-escalation like model-based designs; however, their algorithm remains unchanged throughout the trial, similar to algorithm-based designs, facilitating a simple and effective trial implementation to identifying the optimal MTD. For applications of interval designs, see Ivanova et al., (2007), Ivanova and Kim, (2009), and Ji et al., (2010).
@@
Model-assisted designs identify the MTD by estimating DLT probabilities through the isotonic regression with post-trial data on the number of patients treated and the number of observed DLT events at each dose level. The pool-adjacent violators algorithm (PAVA; Barlow et al.,, 1972; Leeuw et al.,, 2009) is utilized to estimate the DLT probabilities with weighted means based on the variance of the estimates, ensuring compliance with the natural monotonicity assumption in Phase I oncology trials that toxicity increases with dose. Since the isotonic regression independently estimates the DLT probability for each dose level, estimates tend to be unstable in Phase I oncology trials due to the typically small sample sizes. Additionally, the variances of the estimates used as weights for monotonicity in PAVA are similarly affected.
@@
To overcome the limitations of DLT probability estimation using the isotonic regression with PAVA, we propose a novel MTD identification strategy in model-assisted designs that leverages a Bayesian dose-response model, specifically a binomial regression with the number of observed DLTs as the response variable and the dose level as the explanatory variable. By using logit, log-log, or complementary log-log (clog-log) link functions, DLT probability estimation naturally accounts for the monotonicity between dose and toxicity, unlike PAVA, which requires ad hoc adjustments to the estimates. Additionally, the dose-response model allows for the estimation of DLT probabilities at dose levels with small sample size using information from dose levels with larger allocations, thereby enhancing the stability of MTD identification relative to conventional approaches. An additional advantage is that the estimation process considers not only the number of observed DLTs and patients allocated to each dose level but also the dose levels themselves. It is important to note that our approach improves the estimation of DLT probabilities in model-assisted designs using post-trial toxicity data to enhance MTD identification, while preserving the simplicity of trial implementation, unlike model-based designs.
In this study, we focus on the BOIN design, one of the most prominent model-assisted designs, and conduct extensive simulations to examine the impact of using the dose-response model instead of the isotonic regression on MTD selection. Our emphasis on the BOIN design stems from its superior accuracy in identifying the optimal MTD and its lower risk of overdosing compared to the mTPI design, as well as its simpler implementation and ease of execution based on estimated DLT probabilities compared to the Keyboard design (Zhou et al.,, 2018).
@@ 2 Bayesian Optimal Interval design
The BOIN design is one of the interval designs, where dose escalation and de-escalation are determined based on a simple comparison of the observed DLT rate at the current dose level with a predefined toxicity tolerance interval (Oron et al.,, 2011). One of the key advantages of interval designs is that they do not require complex calculations to determine the dose for the next cohort during the trial. The BOIN design consists of three key steps. In the following, we provide a brief explanation of each step.
@@ 2.1 Preparation for trial
Under the Bayesian framework, letting $\pi_{ij}=P(H_{ij})$ $(i=0,1,2)$ denote the prior probability that each hypothesis is true, the probability of making an incorrect decision at each of the dose assignments is give by
@@
$\phi-\Delta_{L}=\frac{\log(\frac{1-\phi_{1}}{1-\phi})}{\log\left(\frac{\phi(1-\phi_{1})}{\phi_{1}(1-\phi)}\right)},\qquad\phi+\Delta_{U}=\frac{\log(\frac{1-\phi}{1-\phi_{2}})}{\log\left(\frac{\phi_{2}(1-\phi)}{\phi(1-\phi_{2})}\right)}$
with the non-informative prior $\pi_{0j}=\pi_{1j}=\pi_{2j}=1/3$. For the detail of the derivation, see Liu and Yuan, (2015) and Yuan et al., (2023). These bounds are independent of the dose level $j$ and the number of subjects $n_{j}$. The specification of $\phi_{1}$ and $\phi_{2}$ has been extensively examined by and it is recommended to set $(\phi_{1},\phi_{2})=(0.6\phi,1.4\phi)$. Therefore, the optimal toxicity tolerance interval for dose escalation and de-escalation can be prespecified during trial planning only by setting the target DLT probability $\phi$.
@@ 2.2 Collected toxicity data
Upper numbers of DLTs $(\phi-\Delta_{U})$ 2 3 4 5 $\cdots$ 12 13
@@ 2.3 Estimation of DLT probability and MTD selection
Let $m_{j}$ denote the total number of DLTs and allocated patients at dose $d_{j}$. In the BOIN design, the DLT probability is estimated using the isotonic regression with the PAVA (Leeuw et al.,, 2009). The procedure is as follows. First, for all dose levels, we calculate
@@ 3.2 Prior distribution for dose-response model
The specification of prior distributions is crucial for incorporating available information about the dose-toxicity relationship, or reflecting the absence of such information. Particularly in cases with limited sample sizes, leveraging reliable external information can enhance estimation accuracy. In the absence of prior information, flat prior distributions (e.g., uniform distributions) or those with large variance are typically used. Assigning such priors to the parameters in the dose-response model (4) leads to a U-shaped prior for DLT probabilities, with spikes at zero and one. While this may be appropriate for the highest and lowest doses, it may result in a misspecified prior for other dose levels, reducing the accuracy of DLT probability estimation. To avoid this issue, following Neuenschwander et al., (2008), we consider a specification of prior distributions that account for the presence or absence of prior information at each dose level.
First, we set $K$ quantiles for the prior distribution of the DLT probability at each dose level $j$, that is, for $k=1,\ldots,K$,
$q_{d_{j}}=\{q_{d_{j}}(p_{1}),\dots,q_{d_{j}}(p_{K})\},\quad P\{\pi({d_{j}})\leq q_{d_{j}}(p_{k})\}=p_{k}.$
Therefore, the prior information is summarized by $J\times K$ quantiles, $Q=(q_{jk})$ with $q_{jk}=q_{d_{j}}(p_{k})$ ($j=1,\ldots,J;k=1,\ldots,K$).
Our prior information on the DLT probability is incorporated through the prior distributions of the coefficient parameters $\beta_{0}$ and $\beta_{1}$ in the dose-response model (4). Consequently, we turn our attention to the specification of the prior distributions for $\beta_{0}$ and $\beta_{1}$. In this paper, we assume $\beta_{0}\sim N(\gamma_{0},\sigma_{0}^{2})$ and $\beta_{1}\sim N(\gamma_{1},\sigma_{1}^{2})$ as prior distributions and discuss the specification of the prior hyperparameters $\eta=\{\gamma_{0},\gamma_{1},\sigma_{0}^{2},\sigma_{1}^{2}\}$. The optimal value of hyperparameter $\eta$ is selected by minimizing the sum of squared errors
$C(Q,Q^{\prime})=\sum_{j,k}(q_{jk}-q^{\prime}_{jk})^{2},$ (5)
@@
where $Q^{\prime}=(q^{\prime}_{jk})$ with the quantile $q^{\prime}_{jk}$ of the DLT probability numerically derived from the prior distributions of $\beta_{0}$ and $\beta_{1}$. While Neuenschwander et al., (2008) employ the loss function $C(Q,\>Q^{\prime})=\max_{j,\>k}|q_{jk}-q^{\prime}_{jk}|$, it suffers from the convergence issues caused by minimizing the maximum absolute value using the stochastic optimization based on the Metropolis algorithm (Robert and Casella,, 2004). The differentiability of the loss function (5) enhances optimization efficiency while allowing the evaluation of errors across all quantiles, leading us to select it as our preferred loss function.
If no prior information is available for the quantiles $Q$, we employ the minimally informative unimodal Beta distribution (Neuenschwander et al.,, 2008) defined as follows.
Definition 1 (Neuenschwander et al.,, 2008).
Given a pre-specified quantile $q(p)$ of the prior distribution, $\pi\sim\text{Beta}(a,b)$ is minimally informative unimodal if (i) $P\{\pi<q(p)\}=p$, (ii) $a\geq 1$ or $b\geq 1$ (or both) and (iii) $a+b$ minimal.
When $q(p)>p$, the hyperparameter $a$ of the minimally informative unimodal Beta distribution $\text{Beta}(a,1)$ is given by $a=\log(p)/\log\{q(p)\}$. Alternatively, when $q(p)<p$, the hyperparameter $b$ of the minimally informative unimodal Beta distribution $\text{Beta}(1,b)$ is given by $b=\log(1-p)/\log\{1-q(p)\}$. We follow the steps below to obtain the prior quantiles $Q$ for the DLT probability using the minimally informative unimodal Beta distribution and incorporate them into the prior distributions of $\beta_{0}$ and $\beta_{1}$.
@@
    Compute the prior median $\mu_{j}=g^{-1}\{\beta^{\prime}_{0}+\exp(\beta^{\prime}_{1})\log(d_{1}/d^{*})\}$ for $j=2,\ldots,J-1$ and derive the hyperparameter of the minimally informative unimodal Beta prior.
@@
Since the BOIN design defines its stopping criterion as the posterior probability of exceeding the target DLT probability at the lowest dose surpassing 95%, we set $P\{\pi(d_{1})>0.3(=\phi)\}=0.05$ in Step 1. Likewise, at $d_{J}$, we assign $P\{\pi(d_{J})\leq 0.1\}=0.05$. These values should be discussed with practitioners when designing trials.
@@ 4.2 Prior distribution of parameters and posterior computation
$\beta_{0}\sim N(-1.592,1.371)\quad\beta_{1}\sim N(0.412,0.784),$
$\beta_{0}\sim N(-0.231,0.847)\quad\beta_{1}\sim N(0.068,0.544),$
@@
$\beta_{0}\sim N(-1.549,0.943)\quad\beta_{1}\sim N(0.142,0.743),$

### v2

# Improving Maximum Tolerated Dose Selection in Model-Assisted Designs for Phase I Trials through Bayesian Dose-Response Model

@@ 1 Introduction
Given these limitations, model-assisted designs (Zhou et al.,, 2018; Yuan et al.,, 2019), which merge the operational simplicity of algorithm-based designs with the good performance of model-based approaches, have attracted growing interest in recent years. Model-assisted designs identify the MTD by estimating DLT probabilities through the isotonic regression with post-trial data on the number of patients treated and the number of observed DLT events at each dose level. The pool-adjacent violators algorithm (PAVA) (Barlow et al.,, 1972; Leeuw et al.,, 2009) is utilized to estimate the DLT probabilities with weighted means based on the variance of the estimates, ensuring compliance with the natural monotonicity assumption in Phase I oncology trials that toxicity increases with dose. Since the isotonic regression independently estimates the DLT probability for each dose level, estimates tend to be unstable in Phase I oncology trials due to the typically small sample sizes. Additionally, the variances of the estimates used as weights for monotonicity in PAVA are similarly affected.
@@
To overcome the limitations of DLT probability estimation using the isotonic regression with PAVA, we propose a novel MTD identification strategy in model-assisted designs that leverages a Bayesian dose-response model, specifically a binomial regression with the number of observed DLTs as the response variable and the dose level as the explanatory variable. By using logit, log-log, or complementary log-log (clog-log) link functions, DLT probability estimation naturally accounts for the monotonicity between dose and toxicity. This contrasts with PAVA, which requires ad hoc adjustments to the estimates. Additionally, the dose-response model allows for the estimation of DLT probabilities at dose levels with small sample size using information from dose levels with larger allocations, thereby enhancing the stability of MTD identification relative to conventional approaches. An additional advantage is that the estimation process considers not only the number of observed DLTs and patients allocated to each dose level but also the dose levels themselves. It is important to note that our approach improves the estimation of DLT probabilities in model-assisted designs using post-trial toxicity data to enhance MTD identification, while preserving the simplicity of trial implementation, unlike model-based designs.
In this study, we focus on the Bayesian optimal interval (BOIN) design (Liu and Yuan,, 2015), one of the most prominent model-assisted designs. BOIN is also one of the interval designs (Ivanova et al.,, 2007; Ivanova and Kim,, 2009; Ji et al.,, 2010), where dose escalation and de-escalation are determined based on the observed DLT rate at the current dose, compared to a predefined toxicity tolerance interval to decide the appropriate dose for the next patient cohort. Oron et al., (2011) demonstrate that interval designs offer convergence properties similar to those of the CRM. We conduct extensive simulations to examine the impact of using the dose-response model instead of the isotonic regression on MTD selection. Our emphasis on the BOIN design stems from its superior accuracy in identifying the optimal MTD and its lower risk of overdosing compared to the mTPI design, as well as its simpler implementation and ease of execution based on estimated DLT probabilities compared to the Keyboard design (Zhou et al.,, 2018). The BOIN framework has been extensively developed to address various clinical scenarios, with iBOIN (Zhou et al.,, 2021) (BOIN with informative prior) allowing incorporation of historical toxicity information through informative priors, gBOIN (Mu et al.,, 2019) (generalized BOIN) generalizing the design to handle continuous, quasi-binary, and binary toxicity endpoints, U-BOIN (Zhou et al.,, 2019) (utility-based BOIN) incorporating utility functions for dose optimization based on risk-benefit trade-offs, and BOIN12 (Lin et al.,, 2020) (BOIN phase I/II) providing one-stage utility-based design for targeted therapies and immunotherapies (see Yuan et al., (2023) for comprehensive reviews of these BOIN variants). These developments demonstrate the adaptability of the model-assisted approach. However, regardless of these extensions, the core MTD selection process in BOIN designs relies on isotonic regression for final dose recommendation. Our proposed modification addresses this fundamental aspect by incorporating parametric dose-response modeling to potentially improve MTD selection accuracy across all BOIN variants.
@@ 2 Bayesian Optimal Interval design
The BOIN design (Liu and Yuan,, 2015) is one of the interval designs, where dose escalation and de-escalation are determined based on a simple comparison of the observed DLT rate at the current dose level with a predefined toxicity tolerance interval (Oron et al.,, 2011). One of the key advantages of the BOIN design is that they do not require complex calculations to determine the dose for the next cohort during the trial. The BOIN design consists of three key steps. In the following, we provide a brief explanation of each step.
@@ 2.1 Preparation for trial
Under the Bayesian framework, letting $\pi_{ij}=P(H_{ij})$ $(i=0,1,2)$ denote the prior probability that each hypothesis is true, the probability of making an incorrect decision at each of the dose assignments is given by
@@
$\phi-\Delta_{L}=\frac{\log\left(\frac{1-\phi_{1}}{1-\phi}\right)}{\log\left(\frac{\phi(1-\phi_{1})}{\phi_{1}(1-\phi)}\right)},\qquad\phi+\Delta_{U}=\frac{\log\left(\frac{1-\phi}{1-\phi_{2}}\right)}{\log\left(\frac{\phi_{2}(1-\phi)}{\phi(1-\phi_{2})}\right)}$
with the non-informative prior $\pi_{0j}=\pi_{1j}=\pi_{2j}=1/3$. For detailed derivations, see the original BOIN paper or the author’s book (Liu and Yuan,, 2015; Yuan et al.,, 2023). These bounds are independent of the dose level $j$ and the number of subjects $n_{j}$. The specification of $\phi_{1}$ and $\phi_{2}$ has been extensively examined and it is recommended to set $(\phi_{1},\phi_{2})=(0.6\phi,1.4\phi)$ (Liu and Yuan,, 2015; Yuan et al.,, 2023). Therefore, the optimal toxicity tolerance interval for dose escalation and de-escalation can be prespecified during trial planning only by setting the target DLT probability $\phi$.
@@ 2.2 Collected toxicity data
Upper numbers of DLTs $(\phi+\Delta_{U})$ 2 3 4 5 $\cdots$ 12 13
@@ 2.3 Estimation of DLT probability and MTD selection
Let $m_{j}$ denote the total number of DLTs and allocated patients at dose $d_{j}$. In the BOIN design, the DLT probability is estimated using the isotonic regression with the PAVA (Barlow et al.,, 1972; Leeuw et al.,, 2009). The procedure is as follows. First, for all dose levels, we calculate
@@ 3.2 Prior distribution for dose-response model
The specification of prior distributions is crucial for incorporating available information about the dose-toxicity relationship, or reflecting the absence of such information. Particularly in cases with limited sample sizes, leveraging reliable external information can enhance estimation accuracy. In the absence of prior information, flat prior distributions or those with large variance are typically used. Assigning such priors to the parameters in the dose-response model (4) leads to a U-shaped prior for DLT probabilities, with spikes at zero and one. While this may be appropriate for the highest and lowest doses, it may result in a misspecified prior for intermediate dose levels, where moderate toxicity rates are typically expected rather than extreme values near zero or one. This prior-data mismatch can lead to biased estimation, which is particularly problematic given the limited sample sizes characteristic of phase I trials. To avoid this issue, following Neuenschwander et al., (2008), we consider approximating the prior distribution for DLT probabilities obtained through the dose-response model (4) with dose-specific prior distributions for DLT probabilities at each dose level. Specifically, we first set quantile targets for DLT probabilities at each dose based on clinical knowledge or design constraints, then derive the hyperparameters of the coefficient priors by minimizing the discrepancy between these target quantiles and the quantiles implied by the dose-response model.
Let $q_{d_{j}}=\{q_{d_{j}}(p_{1}),\dots,q_{d_{j}}(p_{K})\}$ denote the set of $K$ quantiles for the dose-specific prior distribution of the DLT probability $\pi(d_{j})$ at dose level $j$, such that $P\{\pi(d_{j})\leq q_{d_{j}}(p_{k})\}=p_{k}$ for $k=1,\ldots,K$. Our prior information is thus summarized by $J\times K$ quantiles, $Q=(q_{jk})$ where $q_{jk}=q_{d_{j}}(p_{k})$ for $j=1,\ldots,J$ and $k=1,\ldots,K$. On the other hand, let $Q^{\prime}=(q^{\prime}_{jk})$ represent the corresponding quantiles numerically computed from the dose-response model (4) using the prior distributions $\beta_{0}\sim N(\gamma_{0},\sigma_{0}^{2})$ and $\beta_{1}\sim N(\gamma_{1},\sigma_{1}^{2})$. That is, each element $q^{\prime}_{jk}$ represents the $k$-th quantile of $\pi(d_{j})=g^{-1}\{\beta_{0}+\exp(\beta_{1})\log(d_{j}/d^{*})\}$ under the assumed priors for coefficients. Our aim is to determine hyperparameters $\eta=\{\gamma_{0},\gamma_{1},\sigma_{0}^{2},\sigma_{1}^{2}\}$ of the coefficient priors that most faithfully reproduce the dose-specific prior information. To achieve this, we minimize the discrepancy $C(Q,Q^{\prime})$ between the target quantiles $Q$ and the model-implied quantiles $Q^{\prime}$ by solving the optimization problem $\eta^{*}={\rm argmin}_{\eta}C(Q,Q^{\prime})$. We define the loss function as
$C(Q,Q^{\prime})=\sum_{j,k}(q_{jk}-q^{\prime}_{jk})^{2}.$ (5)
@@
While Neuenschwander et al., (2008) employ $C(Q,Q^{\prime})=\max_{j,k}|q_{jk}-q^{\prime}_{jk}|$, which focuses on minimizing the largest quantile error, it suffers from convergence issues caused by minimizing the maximum absolute value using stochastic optimization based on the Metropolis algorithm (Robert and Casella,, 2004). In contrast, our squared error approach considers deviations across all quantiles simultaneously rather than focusing solely on the worst-case error. The differentiability of the loss function (5) enhances optimization efficiency while allowing the evaluation of errors across all quantiles, leading us to select it as our preferred loss function.
In cases where specific prior information about quantiles $Q$ is unavailable, we set the target quantiles $Q$ based on minimally informative unimodal Beta distributions defined by Neuenschwander et al., (2008) The minimally informative unimodal Beta distribution is defined as the Beta distribution $\pi\sim\text{Beta}(a,b)$ that satisfies (i) $P\{\pi<q(p)\}=p$ with the least information ((ii) minimal $a+b$) among all unimodal Beta distributions (where (iii) $a\geq 1$ or $b\geq 1$). A key advantage of this prior distribution is that it requires only a single quantile criterion for each dose level about condition (i) to fully specify the prior distribution. When $q(p)>p$, corresponding to a right-skewed prior, the hyperparameter $a$ of the minimally informative unimodal Beta distribution $\text{Beta}(a,1)$ is given by $a=\log(p)/\log\{q(p)\}$. Conversely, when $q(p)<p$, indicating a left-skewed prior, the hyperparameter $b$ of the minimally informative unimodal Beta distribution $\text{Beta}(1,b)$ is given by $b=\log(1-p)/\log\{1-q(p)\}$.
By leveraging the dose-response model (4), we can further reduce the information required for prior specification. In the following steps, we proceed to obtain the prior quantiles $Q$ for the DLT probability using the minimally informative unimodal Beta distribution and specify the hyperparameters $\eta$ of the coefficient priors to match these target quantiles through the optimization procedure described earlier.
@@
    Compute the prior median $\mu_{j}=g^{-1}\{\beta^{\prime}_{0}+\exp(\beta^{\prime}_{1})\log(d_{j}/d^{*})\}$ for $j=2,\ldots,J-1$ and derive the hyperparameter of the minimally informative unimodal Beta prior.
@@
In Step 2, the complete prior specification across all dose levels can be achieved by considering BOIN’s safety mechanisms and specifying only two probabilities, $p_{1}=P\{\pi(d_{1})>\phi(=0.3)\}$ for the lowest dose and $p_{J}=P\{\pi(d_{J})\leq 0.3\}$ for the highest dose. This approach significantly reduces the burden of prior specification, eliminating the need for detailed prior information for each dose individually. The BOIN design imposes a dose elimination/safety stopping rule. According to this rule, the dose level $j$ and the higher doses are eliminated from the trial if the posterior probability $P\{\pi(d_{j})>0.3|m_{j},n_{j}\}>0.95$ and $n_{j}\geq 3$ (Yan et al.,, 2020; Yuan et al.,, 2023). The trial is stopped if the lowest dose is eliminated. This safety mechanism provides natural constraints for our prior specification, $p_{1}\leq 0.95$ and $1-p_{J}\leq 0.95$, that is, $p_{J}\geq 0.05$. Within these constraints, we address the dual concerns of controlling unacceptable toxicity risk at the lowest dose and avoiding underdosing at the highest dose (Neuenschwander et al.,, 2008). For the lowest dose $d_{1}$, we set $p_{1}=P\{\pi(d_{1})>0.3\}=0.05$, prioritizing control of unacceptable toxicity occurrence. For the highest dose $d_{J}$, we set $p_{J}=P\{\pi(d_{J})\leq 0.3\}=0.21$ yielding $P\{\pi(d_{J})\leq 0.1\}=0.05$, thereby avoiding underdosing. While these values align with BOIN’s safety framework, practitioners should tailor these probability specifications to each trial’s scientific and clinical context, considering regulatory guidance and ethical committee requirements. To examine the impact of different probability choices on MTD selection performance, sensitivity analyses with alternative $p_{1}$ and $p_{J}$ values are provided in Appendix.
@@ 4.2 Prior distribution of parameters and posterior computation
$\beta_{0}\sim N(-1.592,1.371),\quad\beta_{1}\sim N(0.412,0.784),$
$\beta_{0}\sim N(-0.231,0.847),\quad\beta_{1}\sim N(0.068,0.544),$
@@
$\beta_{0}\sim N(-1.549,0.943),\quad\beta_{1}\sim N(0.142,0.743),$
@@ 4.3 Results
Beyond MTD selection accuracy, we also evaluated the safety performance of each method by examining the proportion of overly toxic dose selections (i.e., doses higher than the true MTD). Table 5 shows these results across all scenarios. The log-log link function demonstrated the most conservative behavior, consistently showing the lowest or second-lowest rate of overly toxic dose selections across scenarios 1-8. This conservative characteristic was particularly evident in scenarios 4, 5, 6, and 8. In contrast, the clog-log link function exhibited more aggressive behavior, often resulting in the highest proportion of overly toxic dose selections among the proposed methods. This pattern likely reflects the nature of the clog-log link, which increases toxicity probabilities more gradually at higher doses compared to other link functions. The logit link function struck a balanced performance between the conservative log-log and the more aggressive clog-log approaches. Notably, the conventional PAVA method showed substantially higher rates of overly toxic dose selection in several scenarios (1, 4, and 6), underscoring a key limitation of isotonic regression approaches that rely solely on observed data without structural dose-toxicity modeling. When considering both MTD selection accuracy and safety performance together, the logit model appears to offer the most well-balanced approach among the evaluated methods.
Table 5: The proportion of selecting overly toxic doses (doses higher than the true MTD) across 1,000 simulations in each scenario.
Method PAVA logit log-log clog-log
scenario 1 28.0 24.9 18.2 25.6
scenario 2 0.0 0.0 0.0 0.0
scenario 3 8.8 10.5 5.9 10.9
scenario 4 32.5 35.8 28.3 37.5
scenario 5 16.6 18.1 13.0 19.2
scenario 6 21.5 19.4 13.8 19.9
scenario 7 18.0 24.7 18.6 26.2
scenario 8 14.6 15.8 10.4 17.2